OpenStudy (anonymous):

Evaluate the piecewise defined function at the indicated values. f(x) = {x2 + 4x if x ≤ −1 {x if −1 < x ≤ 1 {−1 if x > 1 f(-4)=..?

OpenStudy (anonymous):

well,first we should say -4 is less\equal than -1 or not.

OpenStudy (anonymous):

and it is! \[-4 \le -1\]

OpenStudy (anonymous):

ok and then..? haha

OpenStudy (anonymous):

i will substitute the -4 to those three equations..? or i will do something else to before substitutuing..?

OpenStudy (anonymous):

it's written for x less/equal than -1 we should use this: x^2 + 4x

OpenStudy (anonymous):

so (-4)^2 + 4(-4) = ?

OpenStudy (anonymous):

16-16...?

OpenStudy (anonymous):

yes! 0

OpenStudy (anonymous):

so do you now what you should do?

OpenStudy (anonymous):

now find f(5) please ;)

OpenStudy (anonymous):

5 > 1 so what we should do?

OpenStudy (anonymous):

f(5) to x^2 + 4x..?

OpenStudy (anonymous):

no! for x < (or equal) -1 we should find x^2 + 4x and 5 is not < -1 , it's >1 . and for x>1 we have -1,so for each x > 1 we have -1! for example: f(2) = -1 f(3) = -1 . . . f(1000) = -1 got it?

OpenStudy (anonymous):

sorry.. didnt get.. it... if x ≤ −1 this should mean x must be less that or equal to -1..? when -4 is subtituted to x^2 + 4x..? i didnt really get this question..im confuse at this questions.. sorry

OpenStudy (anonymous):

well,let's start from the begining. all the functions you have seen are like this: f(x) = x^2 f(x) = x^2 - 2 or etc. here we have another kind of functions that for each x that are less or equal than -1 , we should find x^2 + 4x! so if x≤-1 , then f(x) = x^2 + 4x but if x is bigger than -1 and is less/equal than 1 we just should write x! f(x) = x! for example 0 is bigger than x and less/equal than 1 so f(0) = 0 and for all the x that are more than 1 we have -1...for example 2 is more than z (2>1) so f(2) = -1 , and f(3) = -1 , so for all the Ns that Ns are more than 1, f(N) = -1

OpenStudy (anonymous):

ahh .. wait.. so if x is less than or equal to -1.. f(x) = x^2+4x..? is that it..?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aaahhh ok wait

OpenStudy (anonymous):

there is a given number.. its -4 so those ifs are the clue on which equation you will use to substitute the -4..?

OpenStudy (anonymous):

yes :) we should find where is -4 , is more than 1 or is less/equal than -1 or is between -1 and 1 ? and then use the correct function :)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

hahaha.. FINALLY.. hwahhaha

OpenStudy (anonymous):

got it.. one last question.. how will i read this.. if −1 < x ≤ 1..?

OpenStudy (anonymous):

English is not really my mother-language and i don't know in english countries how do they read this :D but i think we should read this: if x is bigger than -1 and it's less or equal than 1.

OpenStudy (factor):

http://openstudy.com/study#/updates/53c4d627e4b05c273e9d6e44 join now with giveaway

OpenStudy (anonymous):

hwahha.. same my mother-language is not english .. im a filipino.. haha.. maybe thats how it should be read. haha tnx again..

OpenStudy (anonymous):

you're welcome ;)

OpenStudy (factor):

http://openstudy.com/study#/updates/53c4d627e4b05c273e9d6e44 join now with giveaway

OpenStudy (anonymous):

ok @factor :D i'll join.i bumped my practicing studio too.i'll practice LATEX again and then i will join your topic :)

OpenStudy (anonymous):

@factor neeeh no deal i dont know what is latex..? haha

OpenStudy (anonymous):

go to this topic to know : http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A) you can learn how to write LaTeX in the tutorials.

OpenStudy (anonymous):

ok @russeldano if you have given the aswers , you can close your question ;)

OpenStudy (anonymous):

haha.. ok.. haha tnx again.. haha